OpenStudy (anonymous):

Which of the following is the equation of the axis of symmetry of the quadratic equation y = 2x^2 – 12x + 22?

OpenStudy (anonymous):

the formula for the equation of the axis of symmetry is x = -b/2a x = -(-12) / 2(2) x = 12 / 4 x = 3 that is the x coordinate. to find the vertex/turning point, plug in that x value and solve for y. The (x,y) combination is your vertex.

OpenStudy (anonymous):

@livelaughlilz thank you soo much